OpenStudy (anonymous):

Find the GCF. 18x3 and 30x5 A. 90x5 B. 6x3 C. 6x5 D. 90x3

jhonyy9 (jhonyy9):

18^3 = 6*3x^3 30x^5 = 6*5x^3 *x^2 what will be right ?

jhonyy9 (jhonyy9):

common factor what are ?

jhonyy9 (jhonyy9):

6 yes and x^3 yes

jhonyy9 (jhonyy9):

so than will be 6x^3

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

do you understand it ?

OpenStudy (anonymous):

Right, okay. Thank you!

jhonyy9 (jhonyy9):

np was my pleasure